OpenStudy (anonymous):

A parabola has a focus of (-1, 5 4 ) and a directrix of y = 3 4 . What is the parabola’s equation?

OpenStudy (anonymous):

y = ax^2 + bx + c standard form of a parabola

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

can you explain why?

OpenStudy (campbell_st):

this is what you know  well a standard form is (x -h)^2 = 4a(y - k) where (h, k) is the vertex and a is the focal length the distance between the vertex and focus is called the focal length (a) its the same distance from the vertex to the directrix. so you know h = -1 since the line of symmetry is x = -1 from the vertex so 2a = 54 - 34 the k value in the vertex is 54 - a = then substitute h = -1, k = 54 - a and your value for a into the equation above and expand